A particle is projected vertically upward with a velocity of 15 m/s from the point O. Find
the:
(i) maximum height reached.
(ii) time taken for it to return to O.
(iii) time taken for it to be 25 m below O.
1
Expert's answer
2020-06-28T18:51:44-0400

According to Newton's second law


"m{dv\over dt}=F"

By the Law of conservation of energy

"K+P=const"

Let "y" denote the height of the particle above the ground.

Let "v(t)" denote the velocity of the particle at time "t"

"{mv_0^2\over 2}+0={mv^2\over 2}+mgy"

"m{dv\over dt}=-mg"


"v(t)=v_0-gt"

(i) maximum height reached. 


"v=0, y=y_{max}"

"{mv_0^2\over 2}+0=0+mgy_{max}"

"y_{max}={v_0^2\over 2g}"

"y_{max}={(15m/s)^2 \over2(9.8m/s^2)}\approx11.480\ m"

(ii) Let "t_2=" time taken for it to return to "0."  

When the partice reaches maximum height


"v(t_1)=0=v_0-gt_1"

"t_1={v_0\over g}"

"t_2=2t_1={2v_0\over g}"

"t_2={2(15m/s)\over 9.8m/s^2}\approx3.061\ s"

(iii) time taken for it to be 25 m below O.


"v(t)={dy\over dt}"

"y(0)=0, v(t)=v_0-gt"

"y=\int v(t)dt=\int(v_0-gt)dt=y(0)+v_0t-{gt^2\over 2}=""=v_0t-{gt^2\over 2}"

Given "y=-25\ m"


"-25=15t_3-{9.8t_3^2\over 2}, t_3\geq0"

"4.9t_3^2-15t_3-25=0"

"t_3={15\pm\sqrt{(-15)^2-4(4.9)(25)}\over 2(4.9)}="

"={15\pm\sqrt{715}\over 9.8}"

Since "t_3\geq0," we take


"t_3={15+\sqrt{715}\over 9.8}\approx4.259\ s"

The particle will be 25 m below approximately "4.259" seconds after start.